• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Beise

    Log entry time 16:08:13 on June29,2000

    Entry number 45189

    keyword=Shift Summary day shift



    Relatively smooth shift. Change of kinematics to thcm=25. There were
    several minor communications failures, including 3 target IOC reboots
    and 1 CODA reboot.

    8-9 AM Beam Dump problems, use the time to change laser spot and change kinematics to theta_cm = +25
    8:50 AM Target not moving, reboot IOC
    9:15 AM spot move over, target fixed
    9:20 AM 12C in place, ask for 5uA for harp scans scan 948 with coda run 2628
    Note that run header for 2628 still says thcm0. Fixed after.

    10:00 AM continue pointing reference measurements

    10:10AM 4 cm dummy run 2629
    10:15 AM 15 cm dummy run 2630
    10:20 AM Accel IOC crash
    10:40 AM Recover from IOC crash but some accel magnets have to be recycled.
    10:45 AM 10 cm dummy run 2631
    10:53 AM LH2 trigger effi run, 5uA, raster on, all ps=1, run 2632, D.T.=68%
    11:05 AM 2nd trigger eff. run, 10 uA, raster on, run 2633, D.T. = 80%
    11:15 AM PRODUCTION RUN, run 2634, 80uA LH2, collimators IN.
    11:56 AM start run 2635, 80uA LH2, 6msr collimators in
    12:39 PM end run, take out collimators
    12:46 start run 2636
    12:50 Attempt to increase beam current to as high as possible.
    No change in beam current
    13:07 start run 2637
    13:12 lose communication with target IOC, reboot
    13:32 start run 2638
    14:00 start run 2639
    14:25 start run 2640
    14:48 start run 2641
    15:10 end run when ROCs 1,2,3,14 lose communication. End Run failed
    15:15 CODA restarted, run 2642 started
    15:30 lose communication with target IOC, reboot it
    15:42 start run 2643